OpenStudy (anonymous):

Simplify: (1 / x)^-n ...how?

OpenStudy (anonymous):

ok\[\frac{1}{x}=x^{-1}\]so\[(\frac{1}{x})^{-n}=(x^{-1})^{-n}=x^{(-1)(-n)}=x^n\]does that help?

OpenStudy (anonymous):

thanks!